If a two-dimensional cross section of the receiver is modeled by the equation y2 = 24x, and the receiver is 12 in. in diameter, how deep is the parabolic dish? What is the location of the focus?
Answer by lwsshak3(11628)

If a two-dimensional cross section of the receiver is modeled by the equation y2 = 24x, and the receiver is 12 in. in diameter, how deep is the parabolic dish? What is the location of the focus?
***
Equation is that of a parabola that opens right with vertex at origin (0,0)
Its basic form of equation: y^2=4px
For given parabolic dish:
axis of symmetry: y=0
y^2=24x
4p=24
p=6
focus: (6,0) (6 inches to the right of vertex on the axis of symmetry)
..
y-coordinate at top edge of dish=6
x=y^2/24=6^2/24=36/24=3/2
how deep is the parabolic dish? 1.5 inch